OpenStudy (anonymous):

Find the sum or difference. 15/16-7/32 A13/32 B 23/32 C 3/4

OpenStudy (anonymous):

a or c

OpenStudy (anonymous):

Just plug that into a calculator. (If you have one)

OpenStudy (anonymous):

Easy peezy lemon squeezy

OpenStudy (anonymous):

so is it a or c

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

yes what

OpenStudy (anonymous):

a i think

OpenStudy (anonymous):

who aggress with a

OpenStudy (anonymous):

me lol

OpenStudy (anonymous):

a it is

OpenStudy (aripotta):

um i believe it's b

OpenStudy (aripotta):

pretty sure it's b.

OpenStudy (aripotta):

15/16 = 30/32 30/32 - 7/32 = 23/32